# Spectrally Deconfounded Random Forests

### v2

@@ Appendix B Transformation of singular values
The spectral transformation described in Section 3 shrinks the first few singular values to decrease confounding bias. This is visualized in Figure 10 for the PCA adjustment and the trim-transform. PCA adjustment removes all the signal from the first 20 principal components (we assume here that the number of 20 hidden factors is known). This should not reduce the signal of $f^{0}(X)$ as it lies in the span of a sparse set of covariates. However, choosing the right number of principal components in real data is subtle (Owen & Wang 2016) and may result in unwanted removal of signal of $f^{0}(X)$. The trim-transform is much more insensitive to the problem with the unknown number of confounders, as it only limits the top half of the singular values to their median and does not completely remove any principal components and their associated signal completely.
@@ Appendix C Non-linear Confounding
i.e., we are approximately in the setting of Equation 1 with linear confounding with $H$ being replaced by $B$ (and without loss of generality, by orthogonalization, we can assume $\mathrm{Cov}(B)=I$). If $\Gamma$ is dense, it is reasonable that spectral deconfounding still works well. Intuitively, $\Gamma$ will be dense if the $g_{j}$ are all “sufficiently different”. Moreover, $d(\cdot)$ should be “similar” to the $g_{j}$ such that there is an approximation with a common basis. The number $K$ in the basis approximations is then analogous to the number of confounding variables. If we assume that the functions $d(\cdot)$ and $g_{j}(\cdot)$, $j=1,\ldots,p$ are not too complicated, it is still reasonable to assume that $K$ is small and we do not need to know it.
Empirically, we simulate data similar to Section 5. In addition to the non-linear function $f^{0}(X)$, we simulate non-linear confounding using Equation 14 where we simulate $g_{j}$ and $d$ using different random functions using the Fourier basis (also using Equation 13). For the confounding effect on $X$ and $Y$, we use $K=12$ basis functions and sample all the coefficients uniformly on $[-1,1]$ for the effect on $X$ and on $[-2,2]$ for the effect on $Y$. Here, we use $q=1$, $p=300$, $n=500$, and let only one covariate affect the response $Y$. For reasonable noise level and confounding strength, we sample $\delta\in\mathbb{R}$ i.i.d. from a Gaussian with mean zero and $\sigma=2$ and the additional noise $\nu\in\mathbb{R}^{n}$ from a Gaussian with mean zero and $\sigma_{\nu}=0.01$. All the other parameters in the simulation stay the same as in Section 5. In Figure 11, we show the singular values of $\mathbf{X}$ of a random realization. We observe that six instead of just one singular value spike due to the non-linear confounding. Apparently, this number is lower than the number of basis functions $K=12$, and the spiking effect is only visible for the first six components. In this setting, $\mathbf{Y}$ is an even worse approximation for $f^{0}(\mathbf{X})$, but the spectral transformation still results in a clear correlation between $Q\mathbf{Y}$ and $Qf^{0}(\mathbf{X})$, see Figure 12. In Figure 13, we show the dependence of $\mathbf{Y}$, $f^{0}(\mathbf{X})$, and $f^{SDForest}(\mathbf{X})$ on the single causal parent $X_{80}$. The observations (points in the figure) show a complicated dependency to $X_{80}$. At the same time, the estimated relationship by SDForest, represented by the thick line, stays close to the true causal function (dashed line). We repeat this simulation a hundred times and report the test performance, using 500 data points, of classical Random Forests and the SDForest in estimating the true function $f^{0}(\mathbf{X})$. The distribution of these performances is shown in Figure 14, where we clearly see that SDForest outperforms the classical Random Forests in this specific non-linear confounding setting as well.

### v3

@@ Appendix B Transformation of singular values
The spectral transformation described in Section 3 shrinks the first few singular values to decrease confounding bias. This is visualized in Figure 12 for the PCA adjustment and the trim-transform. PCA adjustment removes all the signal from the first 20 principal components (we assume here that the number of 20 hidden factors is known). This should not reduce the signal of $f^{0}(X)$ as it lies in the span of a sparse set of covariates. However, choosing the correct number of principal components in real data is subtle (Owen & Wang 2016) and may result in unwanted removal of signal of $f^{0}(X)$.
The trim-transform is much less sensitive to the problem with the unknown number of confounders, as it only limits the top half of the singular values to their median and does not entirely remove any principal components and their associated signals.
@@ Appendix C Non-linear Confounding
i.e., we are approximately in the setting of Equation 1 with linear confounding with $H$ being replaced by $B$ (and without loss of generality, by orthogonalization, we can assume $\mathrm{Cov}(B)=I$).
If $\Gamma$ is dense, it is reasonable that spectral deconfounding still works well. Intuitively, $\Gamma$ will be dense if the $g_{j}$ are all “sufficiently different”. Moreover, $d(\cdot)$ should be “similar” to the $g_{j}$ such that there is an approximation with a common basis. The number $K$ in the basis approximations is then analogous to the number of confounding variables. If we assume that the functions $d(\cdot)$ and $g_{j}(\cdot)$, $j=1,\ldots,p$ are not too complicated, it is still reasonable to assume that $K$ is small and we do not need to know it.
Empirically, we simulate data similar to Section 5. In addition to the non-linear function $f^{0}(X)$, we simulate non-linear confounding using Equation 14 where we simulate $g_{j}$ and $d$ using different random functions using the Fourier basis (also using Equation 13). For the confounding effect on $X$ and $Y$, we use $K=12$ basis functions and sample all the coefficients uniformly on [1,1][-1,1] for the effect on $X$ and on [2,2][-2,2] for the effect on $Y$.
Here, we use $q=1$, $p=300$, $n=500$, and let only one covariate affect the response $Y$. For reasonable noise level and confounding strength, we sample $\delta\in\mathbb{R}$ i.i.d. from a Gaussian with mean zero and $\sigma=2$ and the additional noise $\nu\in\mathbb{R}^{n}$ from a Gaussian with mean zero and $\sigma_{\nu}=0.01$. All the other parameters in the simulation stay the same as in Section 5.
In Figure 13, we show the singular values of $\mathbf{X}$ of a random realization. We observe that six instead of just one singular value spike due to the non-linear confounding. Apparently, this number is lower than the number of basis functions $K=12$, and the spiking effect is only visible for the first six components.
In this setting, $\mathbf{Y}$ is an even worse approximation for $f^{0}(\mathbf{X})$ compared to the linear confounding setting. However, the spectral transformation still yields a clear correlation between $Q\mathbf{Y}$ and $Qf^{0}(\mathbf{X})$, as shown in Figure 14. In Figure 15, we show the dependence of $\mathbf{Y}$, $f^{0}(\mathbf{X})$, and $f^{SDForest}(\mathbf{X})$ on the single causal parent $X_{80}$. The observations (points in the figure) show a complicated dependency on $X_{80}$.
At the same time, the estimated relationship by SDForest, represented by the thick line, stays close to the actual causal function (dashed line). We repeat this simulation a hundred times and report the test performance, using 500 data points, of classical Random Forests and the SDForest in estimating the actual function $f^{0}(\mathbf{X})$. The distribution of these performances is shown in Figure 16, where we clearly see that SDForest outperforms the classical Random Forests in this specific non-linear confounding setting as well.